# compat=1.0.0

### v1

@@ Abstract
Heavy quark mass in QGP is related to the heavy quark potential at a large distance. In this study we test three different heavy quark potentials, namely, the free energy, the internal energy of the heavy quark pair in QGP, and the unscreened potential, which was recently proposed by the HotQCD Collaboration, through the thermal production of charm quarks in heavy-ion collisions at the LHC. We find that the free energy potential overestimates charm production in heavy-ion collisions at the LHC, while the unscreened potential produces results closest to the experimental data from the ALICE Collaboration among the three potentials.
@@ I Introduction
Some of us have recently studied the thermal production of charm by using the dynamical quasi-particle model Song et al. (2024) which successfully describes the charm scattering in heavy-ion collisions Song et al. (2015, 2016) as well as the spatial diffusion coefficients of heavy quark from the lattice calculations Berrehrah et al. (2014a); Song et al. (2020b). The Feynman diagrams for charm production are simply obtained by rotating those for charm scattering Song et al. (2024). They found that the thermal production overestimates the experimental data on $R_{\rm AA}$ of $D$ meson at LHC, assuming charm quark has the bare mass in QCD Lagrangian, and one possible solution to this is increasing charm quark mass in QGP such that the thermal production is suppressed.
@@
The charm quark mass around $T_{c}$ has also been studied in Ref. Song and Zhou (2024) by matching the charm density or charm fugacity at $T_{c}$ by using the statistical model below $T_{c}$ and the quasi-particle model above $T_{c}$. They have found that charm number density is smoothly connected at $T_{c}$, as expected from the crossover phase transition, if the mass is around 1.8 GeV, which is consistent with the finding in Ref. Song et al. (2024).
@@
On the other hand, the heavy quark mass in QGP is related to heavy quark potential Gubler et al. (2020). Suppose a heavy quark pair is put in QGP. Ignoring kinetic energy, the total energy of the pair will be the sum of masses of heavy quark and heavy antiquark and the potential energy between them. Different from the QED potential, the QCD potential does not converge to zero at infinite distance because of the color confinement. A charm-anticharm quark pair separated from each other infinitely in vacuum will be a charm and anticharm mesons. In QGP, however, the pair will be the dressed charm and anticharm quarks which gain more mass than the bare charm quark.
@@
The study on heavy quark potential is not presently finalized but still in progress. In early days the free energy of heavy quark pair in QGP which is calculable in lattice QCD was suggested as the heavy quark potential Kaczmarek et al. (2004); Burnier et al. (2015). In this case quarknoia melt at relatively low temperatures. On the other hand, the internal energy of heavy quark pair can be considered as the potential Satz (2006). Since the internal energy has the contribution from entropy density, it strongly confines quarknoium near $T_{c}$ and the dissociation temperatures are much higher than those from the free energy potential. One compromise was that the internal energy is more suitable for the initial stage of heavy-ion collisions when a produced heavy quark pair does not have a enough time to exchange energy with the environment, and then the free energy can be used as the potential after a sufficient time elapses. There have also been trials to linearly combine both energies as the heavy potential Wong (2005); Gubler et al. (2020).
@@
Considering the relation between the heavy quark mass in QGP and the heavy quark potential at a long distance, the thermal production of charm quark in heavy-ion collisions will provide useful information about the potential. In this study, we examine the effects of three different heavy quark potentials, the free energy, internal energy and the unscreened potential, on the thermal production of charm and compare with the experimental data in heavy-ion collisions at LHC.
This paper is organized as follows. In Sec. II we briefly review the thermal production of charm quark in QGP and discuss about its dependence on charm quark mass. Then the relation between heavy quark potential and heavy quark mass in QGP is presented. In Sec. III the thermal production is implemented in heavy-ion collisions within the parton-hadron-string dynamics (PHSD), and the results are presented. Finally, a summary is given in Sec. IV.
@@ II Thermal production rate
The dynamical quasi-particle model (DQPM) deals with a strongly interacting partonic matter Berrehrah et al. (2016); Moreau et al. (2019); Soloveva et al. (2020). In the DQPM (anti)quark and gluon gain thermal mass in the spectral form with a pole mass and spectral width, which originate from (anti)quark and gluon complex self-energies in QGP and, as a result, depend on temperature and baryon chemical potential. The complex self-energy is given by the hard thermal loop calculations, but the strong coupling in it is parameterized such that the equation of state from the lattice QCD is reproduced not only at zero but also non-zero chemical potential Moreau et al. (2019); Soloveva et al. (2020).
@@
Figure 2 shows the thermal production rate of charm as a function of temperature for three different charm quark mass. As seen in the figure, the production rate is sensitive to the charm quark mass. For example, if charm quark mass changes from 1.2 GeV to 1.8 GeV, the production rate decreases by the order of 3 at $T_{c}$ and 10 times at $3T_{c}$, even though the mass difference between them is just 0.6 GeV. The reason is that the temperatures are not high enough to produce charm quark pair. The difference will be smaller with increasing temperature.
According to our recent study, the experimental data on $R_{\rm AA}$ of $D$ mesons in Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV favors little thermal production of charm, which is realized by ”massive” thermal charm quark Song et al. (2024). This finding is consistent with the charm quark mass near $T_{c}$, which is obtained from the comparison of the statistical model for temperatures below $T_{c}$ and the quasi-particle model for temperatures above $T_{c}$ Song and Zhou (2024). In fact, this larger mass of charm quark is also consistent with the DQPM, where quark and gluon, which are in principle (almost) massless, gain masses of a few hundred MeV in QGP Moreau et al. (2019); Soloveva et al. (2020).
On the other hand, the heavy quark mass in QGP is related to the heavy quark and antiquark potential, considering that a heavy quark pair infinitely separated from each other will be two open heavy flavors. Different from the Coulomb potential in QED, the QCD potential does not vanish at large distance due to the confinement. When a heavy quark pair is separated by a certain distance, it is more favored to create a light quark pair between them such that the heavy quark pair turns to two open heavy flavors. In QGP the separated heavy quark will not form $D$ mesons but two dressed charm (anti)quarks. In this respect, one can assume
@@
where $m_{0}=1.26$ GeV is charm quark bare mass with which the Schrödinger equation can reproduce charmonium resonances in vacuum Satz (2006); Zhao et al. (2020); Andronic et al. (2024). We note that 1.26 GeV is very close to the charm quark mass in the Particle Data Group Navas et al. (2024). Considering that thermal production rate of charm quark is sensitive to its mass, one can test heavy quark potential through the thermal production in heavy-ion collisions.
There have been several models for the heavy quark potentials. For example, the free energy of heavy quark pair was suggested as the the heavy quark potential Kaczmarek et al. (2004); Kaczmarek and Zantow (2005a). The internal energy which has additional contribution from entropy density is another candidate for the potential Kaczmarek and Zantow (2005b). In the former case the heavy quark potential is relatively weak and quarkonia melt at relatively low temperatures, while the potential from the internal energy is much stronger, especially near $T_{c}$, and quarkonia survive up to high temperatures Satz (2006). The combination of both was also suggested as the heavy quark potential Wong (2005); Gubler et al. (2020). Recently a complex heavy quark potential was extracted from the correlator of temporal Wilson lines at finite temperature in lattice QCD Bazavov et al. (2024). Surprisingly the real part of the heavy quark potential barely changes with increasing temperature, which means no color screening in QGP, and only the imaginary part of potential enhances.
@@
Figure 3: (Upper) charm quark mass as a function of temperature from the free energy and internal energy in comparison with 1.8 GeV and (lower) the corresponding spatial diffusion coefficients $D_{s}$ in comparison with the lattice QCD calculations Banerjee et al. (2012); Altenkort et al. (2024).
Figure 3 shows charm quark mass as a function of temperature in three different heavy quark potentials. We assume that the charm quark mass from Ref. Bazavov et al. (2024) is constantly 1.8 GeV based on the recent study from the statistical model and the quasi-particle model Song and Zhou (2024). One may take $D$ meson mass for the charm quark mass in the third scenario, but it does not change our conclusion qualitatively, as shown in the next section. One finds that charm quark mass from the internal energy is very large near $T_{c}$ because the entropy density $(\partial F/\partial T)$ rapidly changes during the phase transition. It is still larger above $T_{c}$ than that from the free energy due to the non-vanishing contribution from the entropy density. Both free energy and internal energy potentials predict the decreasing charm quark masses with increasing temperature, while the recent lattice QCD study supports a constant charm quark mass Bazavov et al. (2024). However, the spatial diffusion coefficients of charm quark from the three scenarios are not much different from each other and all consistent with the lattice calculations Banerjee et al. (2012); Altenkort et al. (2024), as shown in the lower panel of Fig. 3.
@@ III Thermal charm production in heavy-ion collisions
Figure 4: Thermal production and annihilation of charm quark for three different charm quark masses in 0-10 % central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV.
We display in Fig. 4 the distribution of charm thermal production and annihilation for charm quark mass from the three different heavy quark potentials. Since each parton has a formation time, which is inversely proportional to its energy, thermal production is suppressed at the very early stage in spite of the largest local energy density. The numbers of produced charm quark pairs in mid-rapidity are respectively 6.4, 2.4 and 0.6 for charm quark masses from the free energy, internal energy, and unscreened potentials. On the other hand, the numbers of annihilated charm quark pairs are 0.5, 0.3 and 0.2 on average. Since the free energy potential produces more charm quarks, more annihilation happens due to the larger charm quark number or charm quark density, though the differences are little. Assuming that $d\sigma_{c\bar{c}}^{NN}/dy=1.165$ mb at $\sqrt{s_{\rm NN}}=$ 5.02 TeV Acharya et al. (2019, 2020, 2022a) and average unclear overlap function, $T_{\rm AA}$, in 0-10 % central collisions is 23.3 $\rm mb^{-1}$ Acharya et al. (2020), the expected number of charm pairs produced from the initial hard scattering is roughly 27.1 per rapidity. Simply counting, the thermal production and annihilation will enhance charm by 23 %, 9 % and 2 % for charm quark masses from the free energy, internal energy and unscreened heavy quark potentials. However, considering the shadowing effects, which suppress initial charm production from nucleon-nucleon hard scattering Song et al. (2016); Eskola et al. (2009), the relative contribution from the thermal production will be larger.
@@
Figure 5 shows $R_{\rm AA}$ of $D$ meson as a function of $p_{T}$ in 0-10 % central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV, including thermal charm production for the three different scenarios of charm quark mass, and a comparison with experimental data from the ALICE Collaboration Acharya et al. (2022b). Since thermal charm mostly contributes to the low $p_{T}$ region, it is shown only up to $p_{T}=$12 GeV. In fact, the $R_{\rm AA}$ of $D$ meson above $p_{T}=6$ GeV barely changes by thermal production.
@@
The dashed lines in Fig. 5 show the results from the optimized parameter set. There are 15 parameters in the EPS09, and the colored bands are obtained from two error sets (pset=16,17), which are most sensitive to charm production in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. One can see that charm quark mass from the free energy potential overestimates the experimental data even when including error bars from the shadowing effects. The results from the internal energy is better than those from the free energy, but they still overestimate charm production, because charm quark mass decreases with temperature as shown in Fig. 3, and thermal production is most active at high temperature.
We obtain the best results from the constant charm quark mass, because it suppresses charm thermal production. In this respect the unscreened heavy quark potential in QGP seems consistent with the experimental data on charm production in heavy-ion collisions at LHC.
@@
Since long time ago the thermal production of charm has been anticipated in relativistic heavy-ion collisions especially at LHC. However, the experimental data from LHC doees not show any evidence for the significant thermal production of charm. One possible answer for this is that charm quark gains mass in QGP. This is related to the heavy quark potential in QGP, considering that the energy of heavy quark pair infinitely separated from each other is equivalent to the mass of two dressed heavy quarks.
In this study we have tested three different heavy quark potentials: the free energy, internal energy of heavy quark pair in QGP, and the unscreened potential from the recent lattice QCD study. The cross sections for thermal charm production from quark-antiquark annihilation and two gluon fusion are calculated in the DQPM, and their reverse reactions are also considered for consistency. They are based on the success of heavy quark scattering in the DQPM for the spatial diffusion coefficient of heavy quark from lattice QCD and experimental data on $R_{\rm AA}$ of $D$ meson in heavy-ion collisions. For the simulation of heavy-ion collisions we have used the PHSD, which has been proved successful for the production and dynamics of various kinds of particles, including heavy flavors, in heavy-ion collisions.
Our results show that the free energy potential overestimates charm production, even when considering the uncertainties from the shadowing effects, which are realized by the EPS09 package. The internal energy is better but still overestimates because charm quark mass decreases with temperature, while thermal production mostly occurs at high temperatures. We obtain the most consistent results with the experimental data from the unscreened heavy quark potential in QGP, since charm quark mass does not decrease with temperature.

### v2

@@ Abstract
The effective mass of heavy quark in QGP is related to the heavy quark potential at a large distance. In this study we test different heavy quark potentials, namely, the screened potential such as the free energy and the internal energy of the heavy quark pair in QGP, and the unscreened potential, which was recently proposed by the HotQCD Collaboration, through the thermal production of charm quarks in heavy-ion collisions at the LHC. We find that the free energy potential overestimates charm production in heavy-ion collisions at the LHC, while the unscreened potential produces results closest to the experimental data from the ALICE Collaboration among the three potentials.
@@ I Introduction
Some of us have recently studied the thermal production of charm by using the dynamical quasi-particle model (DQPM) Song et al. (2024) which successfully describes the charm scattering in heavy-ion collisions Song et al. (2015, 2016) as well as the spatial diffusion coefficients of heavy quark from the lattice QCD (lQCD) calculations Berrehrah et al. (2014a); Song et al. (2020b). The Feynman diagrams for charm production are simply obtained by rotating those for charm scattering Song et al. (2024). They found that the thermal production overestimates the experimental data on $R_{\rm AA}$ of $D$ meson at LHC, assuming charm quark has the bare mass in QCD Lagrangian, and one possible solution to this is increasing effective charm quark mass in QGP such that the thermal production is suppressed.
@@
The effective charm quark mass around $T_{c}$ has also been studied in Ref. Song and Zhou (2024) by matching the charm density or charm fugacity at $T_{c}$ by using the statistical model below $T_{c}$ and the quasi-particle model above $T_{c}$. They have found that charm number density is smoothly connected at $T_{c}$, as expected from the crossover phase transition, if the effective mass is around 1.8 GeV, which is consistent with the finding in Refs. Grandchamp et al. (2004); Song et al. (2024).
@@
On the other hand, the effective mass of heavy quark in QGP is related to heavy quark potential Gubler et al. (2020). Suppose a heavy quark pair is put in QGP. Ignoring kinetic energy, the total energy of the pair will be the sum of masses of heavy quark and heavy antiquark and the potential energy between them. Different from the QED potential, the QCD potential does not converge to zero at infinite distance because of the color confinement. A charm-anticharm quark pair separated from each other infinitely in vacuum will be a charm and anticharm mesons. In QGP, however, the pair will be the dressed charm and anticharm quarks which gain more mass than the bare charm quark.
@@
The study on heavy quark potential is not presently finalized but still in progress. In early days the free energy of heavy quark pair in QGP, which is calculable in lattice QCD, was suggested as the heavy quark potential Kaczmarek et al. (2004); Burnier et al. (2015). In this case, quarknoia melt at relatively low temperatures. On the other hand, the internal energy of a heavy quark pair can be considered as the potential Satz (2006). Since the internal energy has the contribution from the entropy density, it strongly confines quarknoium near $T_{c}$, and the dissociation temperatures are much higher than those from the free energy potential. One compromise was that the internal energy is more suitable for the initial stage of heavy-ion collisions, when a produced heavy quark pair does not have enough time to exchange energy with the environment, and then the free energy can be used as the potential after a sufficient time elapses. There have also been trials to linearly combine both energies as the heavy potential Wong (2005); Gubler et al. (2020).
After that, it was found that a Minkowski-time Wilson loop provides the heavy quark potential in the infinite-time limit Laine et al. (2007); Brambilla et al. (2008). Since it is not directly calculable in lattice QCD, spectral functions are first extracted from the Euclidean-time Wilson loop, and then the Minkowski-time Wilson loop is reconstructed from them Rothkopf et al. (2012). However, the extraction of the spectral function has not yet been well established Bala and Datta (2020); Bala et al. (2022).
@@
Considering the relation between the effective heavy quark mass in QGP and the heavy quark potential at a long distance, the thermal production of charm quark in heavy-ion collisions will provide useful information about the potential. In this study, we examine the effects of three different heavy quark potentials, the free energy, internal energy and the unscreened potential, on the thermal production of charm and compare with the experimental data in heavy-ion collisions at LHC.
This paper is organized as follows. In Sec. II we briefly review the thermal production of charm quark in QGP and discuss about its dependence on effective charm quark mass. Then the relation between heavy quark potential and heavy quark mass in QGP is presented. In Sec. III the thermal production is implemented in heavy-ion collisions within the parton-hadron-string dynamics (PHSD), and the results are presented. Finally, a summary is given in Sec. IV.
@@ II Thermal production rate
The dynamical quasi-particle model (DQPM) deals with a strongly interacting partonic matter Berrehrah et al. (2016); Moreau et al. (2019); Soloveva et al. (2020). In the DQPM (anti)quark and gluon gain thermal mass in the spectral form with a pole mass and spectral width, which originate from (anti)quark and gluon complex self-energies in QGP and, as a result, depend on temperature and baryon chemical potential. The complex self-energy is given by the hard thermal loop calculations, but the strong coupling in it is parameterized such that the equation of state from the lattice QCD is reproduced not only at zero but also non-zero chemical potentials Moreau et al. (2019); Soloveva et al. (2020).
@@
Figure 2 shows the thermal production rate of charm as a function of temperature for three different effective charm quark masses. As seen in the figure, the production rate is sensitive to the charm quark mass. For example, if the effective charm quark mass changes from 1.2 GeV to 1.8 GeV, the production rate decreases by three order of magnitude at $T_{c}$ and 10 times at $3T_{c}$, even though the mass difference between them is just 0.6 GeV. The reason is that the temperatures are not high enough to produce charm quark pair. The difference will be smaller with increasing temperature.
Let us briefly explain why we call $m_{c}$ in Fig. 2 the effective charm quark mass. Either in a vacuum or in the thermal medium (below $T_{c}$), the heavy flavor production can be treated by the QCD factorization theory, where the perturbative calculation for the $c\bar{c}$ production is carried out with almost the bare quark mass and the non-perturbative part describes the charm quark to charmed hadrons due to the absence of free quarks in the vacuum. However, this QCD factorization theory fails to describe the low energy scattering. In a vacuum, this low energy scattering is governed by the effective field theories in which the degrees of freedom are charmed hadrons such as the $D$ meson instead of the charm quark.
For example, the threshold energy for charm production in $e^{+}e^{-}$ scattering is not twice charm quark mass, but twice the $D$ meson mass, because two charm quarks must hadronize. If the scattering energy of $e^{+}e^{-}$ is above twice charm quark mass but below twice $D$ meson mass, a produced charm quark pair cannot overcome the string tension which is related to the long-range heavy quark potential and will eventually be annihilated or form a bound state. In this respect the survived charm production is related to the long-range heavy-quark potential, though charm-quark production itself is a hard process occurring on a very short space-time scale.
In this study we focused on the $c\bar{c}$ thermal production in the QGP. In contrast to the vacuum case, the thermally produced charm quarks do not need to be confined in colorless charm hadrons in the QGP. The main question is then how to define the effective charm quark mass in the QGP. One might think that the threshold energy is simply twice the charm-quark bare mass in QGP. Then charm quarks are produced more easily. This concept is analogous to the strangeness enhancement, which was proposed as a signature of QGP formation in early days Rafelski and Muller (1982). However, it is not the case for charm production in heavy-ion collisions at the LHC, where the experimental data on $R_{\rm AA}$ of $D$ mesons in Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV favors little thermal production of charm Song et al. (2024). If the threshold energy were truly twice the bare mass of the charm quark, significantly more charm quarks should have been observed in Pb+Pb collisions at the LHC.
It is well known from thermal field theory that a parton gains thermal mass in the QGP, and the concept of a massive parton is supported by on-shell and off-shell quasi-particle models that reproduce the equation of state of the QGP from lQCD Moreau et al. (2019); Soloveva et al. (2020). Our recent study on the effective charm quark mass near $T_{c}$ is also consistent with it Song and Zhou (2024).
In QGP the separated heavy quark will not form $D$ mesons but two dressed charm (anti)quarks. In this respect, one can assume
@@
where $m_{0}=1.26$ GeV is charm quark bare mass with which the Schrödinger equation can reproduce charmonium resonances in vacuum Satz (2006); Zhao et al. (2020); Andronic et al. (2024). We note that 1.26 GeV is very close to the charm quark mass in the Particle Data Group Navas et al. (2024). In vacuum Eq. (4) is not charm quark mass but $D$ meson mass. In this respect the effective charm quark mass in this study means a half threshold energy for charm production whether in vacuum or in QGP.
Considering that thermal production rate of charm quark is sensitive to its effective mass, one can test heavy quark potential through the thermal production in heavy-ion collisions. Since our calculations need long range potential, it does not matter whether the produced $c\bar{c}$ is the color-singlet or color-octet state, because the long range potential is almost the same Nakamura and Saito (2004). Furthermore, we calculate the cross sections for charm production by using Feynman diagrams which assume long range separation of final state particles (final state particles are represented by plane waves).
There have been several models for the heavy quark potentials. For example, the free energy of heavy quark pair was suggested as the the heavy quark potential Kaczmarek et al. (2004); Kaczmarek and Zantow (2005a). The internal energy which has additional contribution from entropy density is another candidate for the potential Kaczmarek and Zantow (2005b). In the former case the heavy quark potential is relatively weak and quarkonia melt at relatively low temperatures, while the potential from the internal energy is much stronger, especially near $T_{c}$, and quarkonia survive up to high temperatures Satz (2006). The combination of both was also suggested as the heavy quark potential Wong (2005); Gubler et al. (2020). Recently a complex heavy quark potential was extracted from the correlator of temporal Wilson lines at finite temperature in lattice QCD Bazavov et al. (2024). Surprisingly the real part of the heavy quark potential barely changes with increasing temperature, which means no color screening in QGP, and only the imaginary part of potential enhances, though it depends on how to extract the spectral function from the Euclidean-time Wilson loop Bala and Datta (2020); Bala et al. (2022).
@@
Figure 3: (Upper) effective charm quark mass as a function of temperature from the free energy and internal energy in comparison with 1.8 GeV and (lower) the corresponding spatial diffusion coefficients $D_{s}$ in comparison with the lattice QCD calculations Banerjee et al. (2012, 2022); Altenkort et al. (2024).
Figure 3 shows effective charm quark mass as a function of temperature in three different heavy quark potentials. We assume that the effective charm quark mass from Ref. Bazavov et al. (2024) is constantly 1.8 GeV based on the recent study from the statistical model and the quasi-particle model Song and Zhou (2024). One may take $D$ meson mass for the effective charm quark mass in the third scenario, but it does not change our conclusion qualitatively, as shown in the next section. One finds that the effective charm quark mass from the internal energy is very large near $T_{c}$ because the entropy density $(\partial F/\partial T)$ rapidly changes during the phase transition. It is still larger above $T_{c}$ than that from the free energy due to the non-vanishing contribution from the entropy density. Both free energy and internal energy potentials predict the decreasing effective charm quark masses with increasing temperature, while the recent lattice QCD study supports a constant effective charm quark mass Bazavov et al. (2024). However, the spatial diffusion coefficients of charm quark from the three scenarios are not much different from each other and all consistent with the lattice calculations Banerjee et al. (2012); Altenkort et al. (2024), as shown in the lower panel of Fig. 3.
@@ III Thermal charm production in heavy-ion collisions
Figure 4: Thermal production and annihilation of charm quark for three different effective charm quark masses in 0-10 % central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV.
We display in Fig. 4 the distribution of charm thermal production and annihilation for effective charm quark mass from the three different heavy quark potentials. Since each parton has a formation time, which is inversely proportional to its energy, thermal production is suppressed at the very early stage in spite of the largest local energy density. The numbers of produced charm quark pairs in mid-rapidity are respectively 6.4, 2.4 and 0.6 for effective charm quark masses from the free energy, internal energy, and unscreened potentials. On the other hand, the numbers of annihilated charm quark pairs are 0.5, 0.3 and 0.2 on average. Since the free energy potential produces more charm quarks, more annihilation happens due to the larger charm quark number or charm quark density, though the differences are little. Assuming that $d\sigma_{c\bar{c}}^{NN}/dy=1.165$ mb at $\sqrt{s_{\rm NN}}=$ 5.02 TeV Acharya et al. (2019, 2020, 2022a) and average unclear overlap function, $T_{\rm AA}$, in 0-10 % central collisions is 23.3 $\rm mb^{-1}$ Acharya et al. (2020), the expected number of charm pairs produced from the initial hard scattering is roughly 27.1 per rapidity. Simply counting, the thermal production and annihilation will enhance charm by 23 %, 9 % and 2 % for effective charm quark masses from the free energy, internal energy and unscreened heavy quark potentials. However, considering the shadowing effects, which suppress initial charm production from nucleon-nucleon hard scattering Song et al. (2016); Eskola et al. (2009), the relative contribution from the thermal production will be larger.
@@
Figure 5 shows $R_{\rm AA}$ of $D$ meson as a function of $p_{T}$ in 0-10 % central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=$ 5.02 TeV, including thermal charm production for the three different scenarios of effective charm quark mass, and a comparison with experimental data from the ALICE Collaboration Acharya et al. (2022b). Since thermal charm mostly contributes to the low $p_{T}$ region, it is shown only up to $p_{T}=$12 GeV. In fact, the $R_{\rm AA}$ of $D$ meson above $p_{T}=6$ GeV barely changes by thermal production.
@@
The dashed lines in Fig. 5 show the results from the optimized parameter set. There are 15 parameters in the EPS09, and the colored bands are obtained from two error sets (pset=16,17), which are most sensitive to charm production in central Pb+Pb collisions at $\sqrt{s_{\rm NN}}=5.02$ TeV. One can see that effective charm quark mass from the free energy potential overestimates the experimental data even when including error bars from the shadowing effects. The results from the internal energy is better than those from the free energy, but they still overestimate charm production, because effective charm quark mass decreases with temperature as shown in Fig. 3, and thermal production is most active at high temperature.
We obtain the best results from the constant effective charm quark mass, because it suppresses charm thermal production. In this respect the unscreened heavy quark potential in QGP seems consistent with the experimental data on charm production in heavy-ion collisions at LHC.
@@
Since long time ago the thermal production of charm has been anticipated in relativistic heavy-ion collisions especially at LHC. However, the experimental data from LHC doees not show any evidence for the significant thermal production of charm. One possible answer for this is that charm quark gains mass in QGP. This is related to the heavy quark potential in QGP, considering that the energy of heavy quark pair completely separated from each other is equivalent to the mass of two dressed heavy quarks.
In this study we have tested three different heavy quark potentials: the free energy, internal energy of heavy quark pair in QGP, as the screened potentials, and the unscreened potential from the recent lattice QCD study. The cross sections for thermal charm production from quark-antiquark annihilation and two gluon fusion are calculated in the DQPM, and their reverse reactions are also considered for consistency. They are based on the success of heavy quark scattering in the DQPM for the spatial diffusion coefficient of heavy quark from lattice QCD and experimental data on $R_{\rm AA}$ of $D$ meson in heavy-ion collisions. For the simulation of heavy-ion collisions we have used the PHSD, which has been proved successful for the production and dynamics of various kinds of particles, including heavy flavors, in heavy-ion collisions.
Our results show that the free energy potential overestimates charm production, even when considering the uncertainties from the shadowing effects, which are realized by the EPS09 package. The internal energy is better but still overestimates because effective charm quark mass decreases with temperature, while thermal production mostly occurs at high temperatures. We obtain the most consistent results with the experimental data from the unscreened heavy quark potential in QGP, since effective charm quark mass does not decrease with temperature.